OpenStudy (anonymous):

LM has endpoints L(-1,1) and M(-5,-3) find the coordinates of the midpoint of LM ?

OpenStudy (anonymous):

@LieutenantGeneral

OpenStudy (lieutenantgeneral):

-1 + 1 = 0 -5 + -3 = -8 -8 - 0 = -8 -8 = 8 The answer is -8 OR 8

OpenStudy (anonymous):

thanks

OpenStudy (lieutenantgeneral):

ok